Fermi National Laboratory
 

Operation Summary for March 9, 2011 - March 11, 2011

Wednesday March 9

The day shift began with the Fermi accelerator complex in shutdown due to ComEd power system work.

During the day shift, Operations reported that FESS personnel restored power to MI-8 from a scheduled power outage at 10:22 AM; operators performed CAMAC restores. CUB personnel rotated LCW pumps for the Antiproton Source (Pbar) and Linac 55 systems. Water Works personnel discovered a Main Injector (MI) LCW leak on a valve at MI-232 and repaired it.

During the evening shift, operators monitored the New Muon Lab (NML) cryo cavity 2 coupler conditioning. Operators searched and secured Pbar and it’s pre-target area.

Operations turned on all Booster RF systems except for one RF station (BRF5). Operators monitored a Linac RF station (LRF1) and gradually raised its gradient. Operators monitored the Booster vacuum pumps and reset them as needed.

Thursday March 10

The midnight shift began with shutdown in progress.

During the midnight shift, operators searched and secured the Tevatron (TEV) Transfer Hall and F- Sector. Operations reported a problem with the F-Sector interlock system that required Interlock technicians to come in and make repairs. Operators put the TEV separators into conditioning mode. One of the cryo compressors at F0 tripped off; it will be dealt with on Friday. Operations reported problems with the Linac low energy Buncher; this will be looked at during the day shift.

During the day shift, Operations monitored systems. Personnel continued necessary maintenance on all the machines.

During the evening shift, Operators searched and secured MI-8, MI, and the D-Zero Collision Hall. Operators turned on the Pbar power supplies. Linac personnel and operators had beam tuned to the Linac dump by 8 PM. Booster personnel and operators had beam to the Booster dump at 9:05 PM. Operators put the TEV separators into conditioning mode.

Friday March 9

The midnight shift began with the shutdown in progress.

During the midnight shift, operators hipotted MI. Operations reported that the TEV magnet Tilt Meters detected the magnitude 8.9 earthquake off the coast of Japan.

The Plans for Friday and the Weekend

The plans for today are to restore Fermi power to its normal configuration, complete the MI-8 power maintenance, and get back to normal operations with stacking and stashing early this afternoon and then to establish a low luminosity store this evening. The plans for the weekend are to stack, stash, and store.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/11/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory